# Synthetic Text Generation for Training Large Language Models via Gradient Matching

### v1

@@ 5 Experiments
5.1 Experimental settings.
Datasets. We apply our method to different text classification datasets including SST-2 movie reviews (Socher et al., 2013), Tweet emotions (Mohammad et al., 2018), and Rotten tomatoes (Pang & Lee, 2005).
Models. We use the Phi model (Li et al., 2023a) to generate synthetic data and for supervised fine-tuning. In our transfer setting, we fine-tune Llama3 (Dubey et al., 2024) and OPT-1.3B on the generated synthetic texts using Phi.
@@
Baselines. We compare our method with LLM-generated synthetic data with zero-shot and few-shot methods (Li et al., 2023b). We also compare to popular coreset selection methods, namely Herding (Welling, 2009), K-center (Farahani & Hekmatfar, 2009), and Random. Herding greedily selects real samples to match their mean embedding with that of the original data. K-center selects examples to minimize the largest distance between examples that are not selected and their closest selected example.
Hyperparameters. The number of synthetic tokens is set to the average token length of all samples. For the ADMM, the number of updates $T$ is set to 30 and $\rho$ is chosen from the set $\{0.001,0.05,0.01,\ldots,10\}$. To update $\boldsymbol{X}$, we run 50 iterations of Adam with a constant learning rate of 0.008. For the top-$k$ projection, we use $k$ = 200.
@@ 5.2.1 Generating Larger Synthetic Fine-tuning Data in Data-scarce Regime
First, we consider the case where data is scarce for the target task, and we wish to generate a larger synthetic training data based on a small number of examples from the target task. Figure 1 shows the result of applying GrADmm to generate 100 synthetic examples based on only 5, 10, 20, 50 examples randomly selected from the validation data of SST-2, Tweet emotions, and Rotten tomatoes. We see that GrADmm successfully generates high-quality supervised fine-tuning data that can train Phi to a superior performance over that of training on the available validation data. Notably, GrADmm generated synthetic data based on only 5 real examples outperform the real data by 8.6%, 28.5%, and 32.4% on the three datasets. This confirms the effectiveness of GrADmm in the data-scarce regime.
GrADmm outperforms baselines and preserves privacy. Table 1 compares the performance of fine-tuning on synthetic data generated by GrADmm to that of zero-shot and few-shot techniques. It also shows the performance of fine-tuning on subsets of real data selected by herding, K-center, and Random baselines. We note that among all the methods, only the synthetic data generated by GrADmm can preserve the privacy of training data. We see that GrADmm outperforms all the baselines across various datasets and data sizes, by up to 10.4%. Notably, the synthetic data generated by GrADmm has a similar perplexity to that of real data, while having higher perplexity than LLM-generated synthetic data with zero-shot and few-shot methods. This confirms the more diverse nature of the synthetic data generated by GrADmm, compared to LLM generated data.
@@ 5.3 Ablation study
Figure 2: Synthetic examples generated by GrADmm from SST-2.
@@ 6 Conclusion
In this work, we proposed the first theoretically-rigorous method for generating small subsets of synthetic readable text data by matching the gradient of real examples from a target task. We formulated this problem as a discretely constrained non-convex optimization in the embedding space and applied the Alternating Direction Method of Multipliers (ADMM) to iteratively optimizes the embeddings of synthetic examples to match the target gradient, and map them to a sequence of text tokens with low perplexity. We proved that the generated synthetic text can guarantee convergence of the model to a close neighborhood of the solution obtained by fine-tuning on real data. Our extensive experiments on various classification tasks confirmed the effectiveness of our proposed approach.

### v2

@@ 5 Experiments
5.1 Experimental settings
Datasets. We apply GrADmm to different text classification datasets including SST-2 movie reviews (Socher et al., 2013), Tweet emotions (Mohammad et al., 2018), and Rotten tomatoes (Pang & Lee, 2005b).
Model. We use the Phi model (Li et al., 2023a) to generate synthetic data and for supervised fine-tuning.
@@ 5.1 Experimental settings
Baselines. We compare our method with LLM-generated synthetic data with zero-shot and few-shot methods (Li et al., 2023b). We also compare to popular coreset selection methods, namely Herding (Welling, 2009), K-center (Farahani & Hekmatfar, 2009), and Random.
Hyperparameters. The number of synthetic tokens is set to the average token length of all samples. For ADMM, the number of updates $T$ is set to 30 and $\rho$ is chosen from $\{0.001,0.05,0.01,\ldots,10\}$. To update $\boldsymbol{X}$, we run 50 iterations of Adam with lr = 0.008. For the top-$k$ projection, we use $k$ = 200. For DP, we use $\delta$ = 1e-4 and $\varepsilon$ = 0.05.
@@ 5.2 Main results
(a) Last layer gradient difference
(b) $L_{2}$ embedding distance
(c) Effect of $\varepsilon$
Figure 2: Fine-tuning Phi on synthetic texts generated by GrADmm for SST-2. (a) Normalized last-layer gradient error during fine-tuning on synthetic data. Data generated by GrADmm yields significantly smaller gradient errors compared to the zero-shot baseline, indicating closer alignment with real data. (b) $L_{2}$ embedding distance between GrADmm’s synthetic texts to their closest real training data. (c) Effect of $\varepsilon$ on the performance of fine-tuning Phi on GrADmm’s synthetic data.
First, we consider the case where data is scarce for the target task, and we wish to generate a larger synthetic training data based on a small number of examples from the target task. Figure 1 shows the result of applying GrADmm to generate 100 synthetic examples based on only 5, 10, 20, 50 examples randomly selected from the validation data of SST-2, Tweet emotions, and Rotten tomatoes. We see that GrADmm successfully generates high-quality supervised fine-tuning data that can train Phi to a superior performance over that of training on the available validation data. Notably, GrADmm generated synthetic data based on only 5 real examples outperform the real data by 15.7%, 29.7%, and 31.5% on the three datasets. This confirms the effectiveness of GrADmm in the data-scarce regime.
GrADmm outperforms baselines and preserves privacy. Table 1 compares the performance of fine-tuning on synthetic data generated by GrADmm to that of zero-shot and few-shot techniques. It also shows the performance of fine-tuning on subsets of real data selected by herding, K-center, and Random baselines. We note that among all the methods, only the synthetic data generated by GrADmm can preserve the privacy of training data. We see that GrADmm outperforms all the baselines across various datasets and data sizes, by up to 13.1%. Notably, the synthetic data generated by GrADmm has a similar perplexity to that of real data, while having higher perplexity than LLM-generated synthetic data with zero-shot and few-shot methods. This confirms the more diverse nature of the synthetic data generated by GrADmm, compared to LLM generated data.
@@ 5.2.2 Generating Small Synthetic Data Based on Larger Fine-tuning Data
Table 3: SST-2. Divergence (FID) between the (i) training data distribution (Train), (ii) distribution of the few available real examples (Val), (iii) distribution of the 100 GrADmm synthetic data, (iv) distribution of 100 zero-shot synthetic data (Zero).
[…]
5 71.8 44.2 56.0
10 59.8 43.3
20 51.6 39.8
50 40.8 39.7
5.3 Analysis
GrADmm yields similar gradient to real data during fine-tuning. For fine-tune Phi on GrADmm’s synthetic data generated from SST-2, Figure 2(a) illustrates that the normalized last-layer gradient error, i.e. $(\|\nabla_{\boldsymbol{\theta}_{L}}\mathcal{L}(\theta_{t})-\nabla_{\boldsymbol{\theta}_{L}}\mathcal{L}^{s}(\theta_{t})\|)/\|\nabla_{\boldsymbol{\theta}_{L}}\mathcal{L}(\theta_{t})\|$ at the pretrained parameters is small, and this relation holds during fine-tuning. Notably, GrADmm generated data has a much smaller gradient error than the zero-shot baseline during fine-tuning, corroborating its superior performance. Similar results for other datasets and full gradient error can be found in Appendix D.
GrADmm’s synthetic data is close to real data. Table 3 compares (for Figure 1(a)) the embedding divergence (in terms of FID) between the (i) training data distribution, (ii) the distribution of the few available real examples, (iii) the distribution of the 100 synthetic data generated by GrADmm and (iv) the distribution of 100 synthetic data generated using the zero-shot approach. Our synthetic data has a smaller FID, confirming that it has a more similar distribution to that of real training data, compared to baselines. This corroborates the superior performance of GrADmm. While the effectiveness of GrADmm depends on the diversity of the available real examples, our empirical results show that a small number of randomly selected examples can be leveraged to effectively reduce the expected loss.
GrADmm’s synthetic data is yet different from real data. Figure 2(b) shows the histogram of the distances of synthetic examples to their closest real training data. None of the synthetic examples generated by GrADmm are very similar to the real training examples, confirming that our synthetic data is not identical to real examples.
GrADmm preserves privacy. We apply loss-based MIA (Shokri et al., 2017) to the model fine-tuned on GrADmm’s synthetic data generated for SST-2 (Table 1). We select $N=100$ member samples from the training data and $N$ non-member samples. Then, we find the optimal threshold that maximizes the advantages, defined as $2\times(acc-50\%)$, on these $2N$ samples. Finally, we test loss-based MIA with optimal threshold on another $2N$ samples consisting of $N$ members and $N$ non-members. Averaged advantage scores over 10 runs (smaller absolute values indicate better privacy) are $-2.5\%\pm 3.3$ for $\varepsilon=0.05$ and $-2.9\%\pm 5.0$ for $\varepsilon=\infty$. We see that even our non-DP version retains strong privacy, performing only slightly worse than the explicitly differentially private version.
@@ 5.4 Ablation study
Figure 3: Synthetic examples generated by GrADmm from SST-2.
@@
Effect of $\varepsilon$. Figure 2(c) compares the performance of synthetic texts generated with different $\varepsilon$ values. As expected, increasing $\varepsilon$, i.e., relaxing the privacy constraint, generally leads to better performance. Interestingly, when the number of validation samples is limited to just 5, the DP version with $\varepsilon=4$ outperforms the non-DP version ($\varepsilon=\infty$). We hypothesize that the injected noise may enhance generalization during training, consistent with observations from prior work (He et al., 2021).
@@ 6 Conclusion
We proposed the first theoretically-rigorous method for generating privacy-preserving synthetic readable text data by matching the gradient of real examples from a target task. We formulated this problem as a discretely constrained non-convex optimization in the embedding space and applied the Alternating Direction Method of Multipliers (ADMM) to iteratively optimizes the embeddings of synthetic examples to match the noisy target gradient, and map them to a sequence of text tokens with low perplexity. We proved that the generated synthetic text can guarantee convergence of the model to a close neighborhood of the solution obtained by fine-tuning on real data. Our experiments on various classification tasks confirmed the effectiveness of GrADmm.
Impact Statement
This paper proposes a theoretically-rigorous method to generate synthetic text data for fine-tuning LLMs. This can improve the privacy and training efficiency, and be applied to generate synthetic data in scenarios where real data is expensive or hard to obtain. There are many positive potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
This research was partially supported by the National Science Foundation CAREER Award 2146492, CAREER Award 2144985, the NSF-Simons AI Institute for Cosmic Origins, and an Okawa Research Award. We thank Tianjian Huang for their helpful discussion.
@@ C.3.2 Negative Label
(a) IMDB
(b) Sentence polarity
Figure 7: Data-scarce regime. Generating 100 synthetic samples with GrADmm, based on 5, 10, 20, 50 examples from a target task. Synthetic data generated based on only 5 real examples outperforms the real data by 8.9% and 12.5% on the two datasets.
(a) SST-2
(b) Tweet emotions
(c) Rotten tomatoes
Figure 8: Last-layer gradient differences during fine-tuning on synthetic vs. real data. Data generated by GrADmm yields significantly smaller gradient errors compared to the zero-shot baseline, indicating closer alignment with real data.
(a) SST-2
(b) Tweet emotions
(c) Rotten tomatoes
Figure 9: Full gradient differences during fine-tuning on synthetic vs. real data. Data generated by GrADmm yields significantly smaller gradient errors compared to the zero-shot baseline, indicating closer alignment with real data.
(a) Tweet emotions
(b) Rotten tomatoes
Figure 10: $L_{2}$ embedding distance between GrADmm’s synthetic texts to their closest real training data in (left) Tweet emotions and (right) Rotten tomatoes.
Appendix D Additional experiments
Additional datasets. We include two additional datasets namely IMDB (Maas et al., 2011) and Sentence polarity (Pang & Lee, 2005a) in the setting of Figure 1 in Section 5.2.1. Figure 7 shows the result of applying GrADmm to generate 100 synthetic examples based on only 5, 10, 20, 50 examples randomly selected from the validation data of IMDB and Sentence polarity. We see that GrADmm successfully generates high-quality supervised fine-tuning data that can train Phi to a superior performance over that of training on the available validation data. Notably, GrADmm generated synthetic data based on only 5 real examples outperform the real data by 8.9% and 12.5% on the two datasets. This confirms the effectiveness of GrADmm in the data-scarce regime.
Last-layer gradient error. Figure 8 demonstrates the normalized last-layer gradient error with respect to real data, i.e. $(\|\nabla_{\boldsymbol{\theta}_{L}}\mathcal{L}(\theta_{t})-\nabla_{\boldsymbol{\theta}_{L}}\mathcal{L}^{s}(\theta_{t})\|)/\|\nabla_{\boldsymbol{\theta}_{L}}\mathcal{L}(\theta_{t})\|$, remains low at the pretrained parameters at the pretrained parameters and continues to stay low throughout fine-tuning. Notably, the data generated by GrADmm yields a significantly lower gradient error than the zero-shot baseline during fine-tuning, supporting its better performance.
Full gradient error. Figure 9 shows that the normalized full gradient error with respect to real data, i.e. $(\|\nabla_{\boldsymbol{\theta}}\mathcal{L}(\theta_{t})-\nabla_{\boldsymbol{\theta}}\mathcal{L}^{s}(\theta_{t})\|)/\|\nabla_{\boldsymbol{\theta}}\mathcal{L}(\theta_{t})\|$. While GrADmm only matches the last-layer gradients of real and synthetic samples, we observe the same trend as that of last-layer gradient error in Figure 8. This reinforces our last-layer argument in Section 4.3.
Embedding distance. Figure 10 presents a histogram of the distances between synthetic examples and their nearest real training samples. The absence of synthetic examples that are extremely close to real data indicates that GrADmm does not simply replicate real training instances.